Soybean meal is 16% protein; cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320 lb mixture that is 10% protein?
How many pounds of the cornmeal should be in the mixture?

How many pounds of the soybean meal should be in mixture?

Answer by Theo(13342)

soybean is 16% protein.
cornmeal is 8% protein.

x = amount of soybeans
y = amount of cornmeal

x + y = 320

amount of protein in 320 pound mixture if equal to .10 * 320 = 32

amount of protein in x pounds of soybeans is .16 * x
amount of protein in 6 pounds of cornmeal is .08 * y

.16*x + .08*y = .10*320

You have 2 equations that need to be solved simultaneously.

They are:

x + y = 320
.16x + .08y = 32

sole for y in the first equation to get:

y = 320 - x

substitute in the second equation to get:

.16x + .08(320-x) = 32 pounds of protein in the mixture.

simplify to get:

.16x + 25.6 - .08x = 32

combine like terms to get:

.08x + 25.6 = 32

subtract 25.6 from both sides of the equation to get:

.08x = 6.4

divide both sides of the equation by .08 to get:

x = 80

You need 80 pounds of soybeans which means you need 240 pounds of cornmeal.

.16*80 + .08*240 = 12.8 + 19.2 = 32 pounds of protein.

answer is you need 80 pounds of soybeans and 240 pounds of cornmeal.